OpenStudy (sapphiremoon):

Newton's Laws. The only two forces acting on a body have magnitudes of 20N and 35N and their directions differ by 80 degrees. The resulting acceleration has a magnitude of 20 m/s^2. What is the mass of the body? I have a solution but I don't understand it. Could someone go over this with me?

OpenStudy (sapphiremoon):

@IrishBoy123 @jiteshmeghwal9

OpenStudy (irishboy123):

 Step 1: draw the forces Step 2: use either one of the force directions as the x -axis or y-axis, then go to Step 3. [in the drawing, we use the 20N direction as the x-axis ] Step 3: resolve forces in x and y directions Step 4: work out the net force using Pitagoras Theorem Here you should get \(F_{net} = \sqrt { (20 + 35 \cos 80 )^2 + (35 \sin 80)^2 } \) And then whack it with... \(F_{net} = m a\) Ta-Da!! ;-)

OpenStudy (sapphiremoon):

Okay, I think my problem was the way I was drawing my diagram made it difficult to understand the forces (with the idea of gravity clouding my brain). I'll see if I can get it. Thanks!